HUBBLE SPACE TELESCOPE DAILY REPORT #1454 PERIOD COVERED: 0000Z (UTC) 08/14/95 - 0000Z (UTC) 08/15/95 Daily Status Report as of 227/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FGS 6157 (Parallaxes of Hyades Cluster Members, Cycle 5 - High) The FGSs were used to make several POS Mode astrometric observations of members (set VA310, AST50-REF, et al., and set PCEN383, VA383, AST64-REF, et al.) of the Hyades Cluster in order to measure their parallaxes. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking ultraviolet observations running concurrently with prime science observations. The intent is to search for extra-galactic sources with flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The observations were completed as planned, and no problems were reported. 1.3 Completed Three Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: Dark) The WF/PC-2 was used to make internal dark calibration images. The observations were completed as scheduled, and no problems were reported. 1.4 Completed FOS/RD 5825 (The Structure of Triton's Lower Atmosphere) The Faint Object Spectrograph (Red Side) was used to make rapid mode spectrographic observations of Neptune's moon Triton to study its lower atmosphere during an occultation with a 12.9 magnitude star. The observations were completed as planned, and no problems were noted. 1.5 Completed WF/PC-2 5972 (Determination of the Extragalactic Distance Scale) The WF/PC-2 was used to make observations of the galaxy NGC-1365 looking for standard candle Cepheid variable stars. The purpose of this observations is to help estimate the Hubble constant with an accuracy of 10%. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 6190 (WF/PC-2 Cycle 4 Calibration: Internal Flats) The WF/PC-2 was used to make an internal flat filter sweep and an internal preflash level test. The internal lamp is also used as a source for the preflash test. The observation was executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration images. Both photometric and narrow-band filters were used. The observations were executed as planned, and no problems were reported. 1.8 Completed WF/PC-2 6129 (The Structure, Polarization, and Evolution of the Crab Synchrotron Nebula) The WF/PC-2 was used to make observations of the area near Crab-Pulsar (M-1) in order to study its structure, polarization, and evolution. Two color and polarization images were made. The observations were executed as planned, and no anomalies were reported. 1.9 Completed Two Sets of FGS 6158 (Parallaxes of Hyades Cluster Members, Cycle 5 - Part 2) The FGSs were used to make astrometric POS Mode observations (Part Two) of members (set VA627, AST113-REF, VA622, et. al. and set PCEN548, VA548, AST96 REF, et. al.) of the Hyades Cluster in order to measure their parallaxes. This is part two of proposal #6157. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of WF/P-2 6254 (GTO Pure Parallel Program - Cycle 5) The WF/PC-2 was use to make random long exposures of high galactic fields taken in parallel with the prime science observations to obtain a random sample of galaxies. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 7 Successful : 7 2.3 Operations Notes: For one scheduled TDRSS support period no data was received because the transmitter command had been omitted from the command load. There was no impact since the ETR was on for the entire support period. A TTR was written to document this event. Several ground system problems were encountered. The PRS3 system failed, causing a work order to be filed. Later the link between the TAC2 and AP1 dropped out temporarily. Still later, AP1 stopped processing data. A system reboot cleared this problem. The SMGT-21 dry-run will be conducted today. The formal run will occur on Friday. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-